Translate the problem into a pair of linear equations in two vaiables. Solve the equations using either elimination or substitution.
A plane flies 400 miles with the wind and 300 miles against the wind in the same length of time. If the speed of the wind is 22 mph, what is the speed of the plane in still air?
I don't know where to begin with this problem.... Please Help.

Answer by mananth(12270)

A plane flies 400 miles with the wind and 300 miles against the wind in the same length of time. If the speed of the wind is 22 mph, what is the speed of the plane in still air?
let speed in still air be xmph
..
speed with wind will be x+22 mph it goes 400 miles
speed against wind will be x-22 . it goes 300 miles.
time taken is the same
..
400/ x+22 = 300/ x-22
400*(x-22)=300(x+22)
400x-8800=300x-6600
100x=15000
x=150 mph the speed of plane in still air